A 3x10^-6 C Negative charge produces a force of 2X10^-2 N at a distance of o.ooo4 m with another unknown charge. What is the magnitude of the unknown charge?


1
Expert's answer
2021-10-10T16:04:22-0400

electrostatic force:

"F=k\frac{q_1q_2}{r^2}"


"q_2=\frac{Fr^2}{kq_1}=\frac{2\cdot10^{-2}\cdot16\cdot10^{-8}}{9\cdot10^9\cdot3\cdot10^{-6}}=1.185\cdot10^{-13}" C